OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 3x – 4 9x – 3y = 14 A. There is no solution. B. (12, 14) C. (14, 12) D. There are an infinite number of solutions.

OpenStudy (anonymous):

A

OpenStudy (jdoe0001):

\(\bf {\color{red}{ y}} = 3x - 4\\ 9x - 3y = 14 \\ \quad \\ 9x - 3y = 14\implies 9x - 3({\color{red}{ 3x - 4}}) = 14\) solve for "x" once you found "x", substitute on either equation and solve for "y"

OpenStudy (anonymous):

the lines are parallel if you graph it